# Gauge-Invariant Entire-Function Regulators and UV Finiteness in Non-Local Quantum Field Theory

### v3

@@ I Entire-function regulators and spectral calculus
The gauge- and diffeomorphism-covariant derivative $D_{\mu}$ acting on a field $\Phi$ is defined by [^1] There are past publications on nonlocal QFT. We provide a list of references for these publications in [9, 10, 7, 8, 4, 5, 6, 2, 3, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 26, 24, 25]:
@@
where $\nabla^{\text{LC}}_{\mu}$ is the Levi–Civita covariant derivative associated with the metric $g_{\mu\nu}$, and $A_{\mu}$ is the gauge connection one-form in the appropriate representation of $\mathfrak{g}$. The covariant Laplace–Beltrami operator, also referred to as the Bochner d’Alembertian in this context, is then given by:
@@
where $M_{*}$ is a constant mass scale that sets the onset of nonlocality. To understand the action of $F(\Box/M_{*}^{2})$, it is useful to recall the spectral calculus for self-adjoint operators. On a globally hyperbolic spacetime one can define a suitable Hilbert space of fields, and at least in the Euclidean or Riemannian case the analogue of $\Box$ is an essentially self-adjoint operator. We then have a spectral decomposition:
@@ III From Minkowski to Euclidean momentum space
where the overall factor of $i$ has been absorbed into the continuation of the Green’s function to Euclidean signature. Comparing (30) with (35), we see that the factor $e^{-p^{2}/M_{*}^{2}}$ in Minkowski space has become $e^{-p_{E}^{2}/M_{*}^{2}}$ in Euclidean space. The crucial point is that $p_{E}^{2}$ is positive and grows without bound as $|p_{E}^{\mu}|\to\infty$, so the exponential factor provides strong UV damping:
@@ IV Gauge invariance and the background-field method
So far we have considered only a scalar field with no gauge interactions. In HUFT and in more general gauge theories, we require the regulator to respect gauge invariance and BRST symmetry. This is achieved by constructing the regulator as an entire function of the covariant operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, rather than of the ordinary Laplacian.
@@ V Liouville’s theorem and the singularity at infinity
We address the concern raised by Liouville’s theorem. Liouville’s theorem states that any bounded entire function must be constant, so any nonconstant entire function must be unbounded and has an essential singularity at the point $z=\infty$ on the Riemann sphere [27, 28].
In our context, $z$ is effectively identified with the complexified eigenvalue of $\Box/M_{*}^{2}$, which in the perturbative vacuum is $p^{2}/M_{*}^{2}$ in Minkowski space or $p_{E}^{2}/M_{*}^{2}$ in Euclidean space. The fact that $F$ is unbounded as $|z|\to\infty$ simply reflects the fact that $F$ must grow in certain directions in the complex plane. However, the physical amplitudes are constructed from the values of $F(z)$ along specific directions such as the real Euclidean axis $z=-p_{E}^{2}/M_{*}^{2}\leq 0$, where $p_{E}^{2}$ is real and nonnegative, the real Minkowski $p^{2}$-axis, approached with the standard $i\epsilon$-prescription for timelike momenta. By choosing $F(z)$ to be free of zeros in the finite complex plane and to have suitable decay along the Euclidean axis, we guarantee that the propagators do not acquire additional poles or branch cuts in the finite $p^{2}$-plane; the only singularities correspond to the physical particle masses. Loop integrals are exponentially damped along the Euclidean axis, yielding UV-finite amplitudes.
@@ VI Nonlocality, microcausality, and the local limit
and locality is recovered. For finite but large $M_{*}$, the theory is quasi-local, with deviations from strict microcausality confined to scales $\lesssim\ell_{*}$. In HUFT, $M_{*}$ is taken to be near the Planck or GUT scale, so these effects are negligible at accessible energies and distances.
@@ VII Paley–Wiener bounds and quasi-locality
The relation between entire-function regulators in momentum space and quasi-locality in position space can be formulated more precisely in terms of Paley–Wiener–type theorems [29, 30, 31]. In flat space, acting on a scalar field, we have:
@@
Given HUFT’s structural assumptions of an entire function regulator with a proper-time representation and the Paley–Wiener–type locality bounds, the admissible regulators form a Paley–Wiener class of entire form factors that exhibit Gaussian-type exponential damping at large Euclidean momentum, up to subleading exponentially suppressed corrections and redefinitions of the nonlocality scale $M_{\ast}$. Our choice $F(z)=e^{z}$ is the simplest representative of this class. From the Paley–Wiener theorem, we learn that the regulator must lie in the class of Gaussian entire functions with asymptotic behaviour of $e^{-}p^{2}_{E}/M_{*}^{2}$, up to subleading polynomial factors and redefinitions of M*
VIII Discussion and outlook

### v5

# On Gauge-Invariant Entire-Function Regulators and UV Finiteness in Non-Local Quantum Field Theory

@@ I Entire-function regulators and spectral calculus
The gauge- and diffeomorphism-covariant derivative $D_{\mu}$ acting on a field $\Phi$ is defined by  MT:HUFT-EPJC ; MT:Invariant ; Moffat:FiniteNonlocal1990 ; MT:FiniteHolomorphicQFT ; MT:SMmass ; MT:SL2C ; MT:ReplyToCline ; Moffat:UVcompleteQG2011 ; Moffat:QGCCP2014 ; Tomboulis1997 ; Biswas2012 ; Talaganis2015 ; Buoninfante2018 ; LM:2023 ; EMKW:1991 ; GE:1967 ; KW:1992 ; M:2011a ; M:2019 ; M:2021 ; GM:2021 ; M:2011b ; AE:1973 ; PS:1995 ; MJN:2011 :
@@
where $\nabla^{\text{LC}}_{\mu}$ is the Levi–Civita covariant derivative associated with the metric $g_{\mu\nu}$, and $A_{\mu}$ is the gauge connection one-form in the appropriate representation of $\mathfrak{g}$. The covariant Laplace–Beltrami operator, also referred to as the Bochner d’Alembertian in this context, is then given by A.Ibort2012 :
@@
where $M_{*}$ is a constant mass scale that sets the onset of nonlocality. This series is convergent as an ordinary complex series, with radius of convergence at infinity, as an operator series defined via spectral/functional calculus, and in the perturbative vacuum it becomes a pointwise statement in momentum space. To understand the action of $F(\Box/M_{*}^{2})$, it is useful to recall the spectral calculus for self-adjoint operators. On a globally hyperbolic spacetime one can define a suitable Hilbert space of fields, and at least in the Euclidean or Riemannian case the analogue of $\Box$ is an essentially self-adjoint operator. We then have a spectral decomposition:
@@ III From Minkowski to Euclidean momentum space
where the overall factor of $i$ has been absorbed into the continuation of the Green’s function to Euclidean signature. Comparing (30) with (35), we see that the factor $e^{-p^{2}/M_{*}^{2}}$ in Minkowski space has become $e^{-p_{E}^{2}/M_{*}^{2}}$ in Euclidean space. For Gaussian entire regulators such as $F(z)=e^{z}$, the safest and standard interpretation of Wick rotation is the following, we define the regulated amplitude first by the absolutely convergent Euclidean integral, and then define the Lorentzian or Feynman amplitude by analytic continuation in the external energies or invariants with the usual $i\epsilon$ prescription. Whenever a literal contour deformation in $p^{0}$ is justified, it reproduces the same analytic continuation. The crucial point is that $p_{E}^{2}$ is positive and grows without bound as $|p_{E}^{\mu}|\to\infty$, so the exponential factor provides strong UV damping:
@@ IV Gauge invariance and the background-field method
So far we have considered only a scalar field with no gauge interactions. In general gauge theories, we require the regulator to respect gauge invariance and BRST symmetry. This is achieved by constructing the regulator as an entire function of the covariant operator $\Box=-g^{\mu\nu}D_{\mu}D_{\nu}$, rather than of the ordinary Laplacian.
@@ V Liouville’s theorem and the singularity at infinity
We address the concern raised by Liouville’s theorem. Liouville’s theorem states that any bounded entire function must be constant, so any nonconstant entire function must be unbounded and has an essential singularity at the point $z=\infty$ on the Riemann sphere BJ:2024 ; H:2022 . In our context, $z$ is effectively identified with the complexified eigenvalue of $\Box/M_{*}^{2}$, which in the perturbative vacuum is $p^{2}/M_{*}^{2}$ in Minkowski space or $p_{E}^{2}/M_{*}^{2}$ in Euclidean space. The fact that $F$ is unbounded as $|z|\to\infty$ simply reflects the fact that $F$ must grow in certain directions in the complex plane. However, the physical amplitudes are constructed from the values of $F(z)$ along specific directions such as the real Euclidean axis $z=-p_{E}^{2}/M_{*}^{2}\leq 0$, where $p_{E}^{2}$ is real and nonnegative, the real Minkowski $p^{2}$-axis, approached with the standard $i\epsilon$-prescription for timelike momenta. By choosing $F(z)$ to be free of zeros in the finite complex plane and to have suitable decay along the Euclidean axis, we guarantee that the propagators do not acquire additional poles or branch cuts in the finite $p^{2}$-plane; the only singularities correspond to the physical particle masses. Loop integrals are exponentially damped along the Euclidean axis, yielding UV-finite amplitudes.
@@
We assume throughout:
(A1) Entire and zero-free at finite $z$. $F:\mathbb{C}\to\mathbb{C}$ is entire and satisfies $F(z)\neq 0$ for all finite $z\in\mathbb{C}$.
(A2) Euclidean UV damping. There exist constants $C>0$ and $a>0$ such that along the Euclidean ray $z=-s\leq 0$,
$|F(-s)|\;\leq\;C\,e^{-as}\qquad\text{for all sufficiently large }s\geq 0.$ (39)
For $F(z)=e^{z}$ one has $|F(-s)|=e^{-s}$ so (39) holds with $C=1$ and $a=1$. Assumption (A1) is the condition of no additional poles or branch cuts from the regulator, only the usual physical singularities remain. Assumption (A2) is precisely the condition used in the Euclidean discussion around to guarantee strong UV damping.
Lemma 1 (the regulator cannot create finite $p^{2}$-plane singularities). Consider a regulated scalar propagator
$\Delta_{F}^{\rm reg}(p)\;=\;\frac{i\,F\!\left(-p^{2}/M_{*}^{2}\right)}{p^{2}-m^{2}+i\epsilon}.$ (40)
Under (A1), the only singularities of $\Delta_{F}^{\rm reg}(p)$ as a function of $p^{2}$ in the finite complex plane are those of the denominator $p^{2}-m^{2}+i\epsilon$, such as the physical pole at $p^{2}=m^{2}-i\epsilon$ and, in interacting theories, the usual multi-particle branch cuts coming from loop integration, not from $F$.
The map $p^{2}\mapsto-p^{2}/M_{*}^{2}$ is entire, and by (A1) $F$ is holomorphic and nonvanishing at every finite value of its argument. Hence $F(-p^{2}/M_{*}^{2})$ has no poles and no zeros for finite $p^{2}$. Multiplying by such a factor cannot introduce poles or branch points at finite $p^{2}$; therefore the only finite-plane pole is the one already present in the local propagator. The same statement applies diagram-by-diagram, that multiplying propagators and vertices by entire, zero-free factors cannot generate new finite kinematic singularities. The Landau singularities of threshold cuts are determined by pinches of the usual denominators; the regulator modifies only the UV weight of large loop momenta, not the on-shell conditions.
Lemma 2 (absolute convergence of Euclidean loop integrals). Fix a Feynman graph $G$ with $L$ loops in $d=4$ Euclidean dimensions and masses $\{m_{j}\}$. Assume each internal line carries a factor $F(-k_{E}^{2}/M_{*}^{2})$ with $F$ satisfying (A2). Then the Euclidean momentum integral defining the regulated amplitude is absolutely convergent in the UV.
At large Euclidean momentum one has $k_{E}^{2}\sim\Lambda^{2}$ for some scaling parameter $\Lambda\to\infty$. The local unregulated integrand grows at most polynomially in $\Lambda$, power-counting. By (A2) each regulated line contributes a factor bounded by $Ce^{-ak_{E}^{2}/M_{*}^{2}}$, so the full integrand is bounded by a polynomial in $\Lambda$ times $\exp(-c\Lambda^{2})$ for some $c>0$. The integral of a polynomial times $\exp(-c\Lambda^{2})$ over $\mathbb{R}^{4L}$ is finite. Therefore the Euclidean integral converges absolutely in the UV.
Proposition 3 (Euclidean-first definition implies no imprint from $z=\infty$). Define the regulated Euclidean Green function, or amplitude for fixed external Euclidean momenta by the absolutely convergent Euclidean loop integral, and its multi-loop generalizations. Then the corresponding Lorentzian or Feynman Green function is defined by analytic continuation of the external energies or invariants with the standard $i\epsilon$ prescription. Under (A1)–(A2), this continuation that (i) is well-defined with no obstruction from the regulator, introduces no new finite-plane singularities beyond the standard physical ones, and does not depend on the essential singularity of $F$ at $z=\infty$.
By Lemma 2, the Euclidean integral converges absolutely. Moreover, for external momenta restricted to any compact set in a complex neighbourhood that avoids the usual denominator pinches, the Euclidean integrand is jointly analytic in the external parameters and dominated by an integrable function, the same polynomial times Gaussian bound used in Lemma 2. Hence, by standard results, dominated convergence or Morera’s theorem, the Euclidean amplitude is an analytic function of the external invariants in that domain.
The only possible singularities of this analytic function upon continuation arise when the continuation crosses the loci where denominators can pinch, the usual Landau or threshold conditions. By Lemma 1, the regulator cannot create additional finite-plane singularities because it is entire and zero-free at finite argument. Thus the physical singularity structure is unchanged, so the regulator alters UV behaviour, not the location or nature of physical thresholds.
The essential singularity at $z=\infty$ is a statement about the behaviour of $F(z)$ as $|z|\to\infty$ in directions of the complex $z$-plane. In the Euclidean-first construction, the only large-$|z|$ region probed by the integral is the Euclidean ray $z=-p_{E}^{2}/M_{*}^{2}\leq 0$, where (A2) enforces decay, ensuring convergence. Analytic continuation of the resulting finite analytic function is performed in the external invariants; it does not require evaluating $F$ on contours that encircle or approach $z=\infty$ in directions where $F$ may grow. Therefore the essential singularity at $z=\infty$ has no operational imprint on the regulated amplitudes.
The statement that the singularity at $z=\infty$ is never encountered should be understood precisely as follows, that physical amplitudes are defined by Euclidean integrals on the decaying ray $z=-p_{E}^{2}/M_{*}^{2}\leq 0$ and then continued analytically with the Feynman $i\epsilon$ prescription; at no stage does the construction require probing the growth sectors associated with $z=\infty$.
For completeness, one may ask when the Euclidean result can also be obtained by a literal contour rotation in the complex $p^{0}$-plane at fixed spatial momentum. This requires an additional sectorial bound on the regulator ensuring that the contribution from the large arc vanishes. A sufficient condition is that for some $\alpha>0$ and some wedge connecting the real and imaginary axes:
$\left|F\!\left(-\frac{p^{2}}{M_{*}^{2}}\right)\right|\;\leq\;\exp\!\left(-\alpha\,|p^{0}|^{2}/M_{*}^{2}\right),$ (41)
as $|p^{0}|\to\infty$ in that wedge, so that Jordan-type estimates force the arc contribution to vanish. When such a bound holds, the Euclidean-first analytic continuation coincides with an actual contour deformation. In our framework, however, the Euclidean-first definition above is sufficient and does not require assuming (41).
@@ VI Nonlocality, microcausality, and the local limit
and locality is recovered. For finite but large $M_{*}$, the theory is quasi-local, with deviations from strict microcausality confined to scales $\lesssim\ell_{*}$. In our non-local QFT, $M_{*}$ is taken to be near the Planck or GUT scale, so these effects are negligible at accessible energies and distances.
@@ VII Paley–Wiener bounds and quasi-locality
The relation between entire-function regulators in momentum space and quasi-locality in position space can be formulated more precisely in terms of Paley–Wiener–type theorems PaleyWiener1934 ; HormanderPW ; SteinWeiss . In flat space, acting on a scalar field, we have:
@@
Given our non-local QFTs structural assumptions of an entire function regulator with a proper-time representation and the Paley–Wiener–type locality bounds, the admissible regulators form a Paley–Wiener class of entire form factors that exhibit Gaussian-type exponential damping at large Euclidean momentum, up to subleading exponentially suppressed corrections and redefinitions of the nonlocality scale $M_{\ast}$. Our choice $F(z)=e^{z}$ is the simplest representative of this class. From the Paley–Wiener theorem, we learn that the regulator must lie in the class of Gaussian entire functions with asymptotic behaviour of $e^{-}p^{2}_{E}/M_{*}^{2}$, up to subleading polynomial factors and redefinitions of $M_{*}$
VIII Osterwalder–Schrader axioms and Euclidean reconstruction
A technically clean way to define regulated amplitudes is the Euclidean-first prescription where one defines the regulated Euclidean Green functions by absolutely convergent Euclidean integrals and then defines the Lorentzian/Feynman correlators by analytic continuation of the external invariants with the standard $i\epsilon$ prescription. This is precisely the setting in which the Osterwalder–Schrader (OS) axioms provide a rigorous bridge between Euclidean correlation functions, the Schwinger functions and a relativistic quantum field theory in Hilbert space OS73 ; OS75 ; GJ81 .
Foe Euclidean $n$-point Schwinger functions we fix Euclidean space $\mathbb{R}^{4}$ with coordinates:
$x_{E}\;=\;(\tau,\mathbf{x}),\qquad\tau\in\mathbb{R},\ \mathbf{x}\in\mathbb{R}^{3},$ (50)
and Euclidean scalar product $x_{E}\cdot y_{E}:=\tau\sigma+\mathbf{x}\cdot\mathbf{y}$. Let $\Delta_{E}:=\partial_{\tau}^{2}+\nabla^{2}$ denote the positive Euclidean Laplacian, and let $M_{\ast}>0$ be the nonlocality scale. In a regulated scalar model one may write a quadratic Euclidean action in the schematic form:
$S_{E}^{(2)}[\phi]:=$ $\frac{1}{2}\int_{\mathbb{R}^{4}}d^{4}x_{E}\;\phi(x_{E})\,$
$\times\Bigl[F(-\Delta_{E}/M_{\ast}^{2})\,(-\Delta_{E}+m^{2})\Bigr]\phi(x_{E}),$ (51)
where $\phi$ is a real Euclidean scalar field, $m\geq 0$ is a mass parameter, and $F$ is the entire regulator, in momentum space $\Delta_{E}\mapsto p_{E}^{2}$). Interactions are encoded by an additional term $S_{E,\mathrm{int}}[\phi]$.
The Euclidean $n$-point functions are defined formally by:
$S_{n}(x_{E,1},\ldots,x_{E,n})$ $:=\frac{1}{Z_{E}}\int\mathcal{D}\phi\;e^{-S_{E}[\phi]}\,\phi(x_{E,1})\cdots\phi(x_{E,n}),$
$\text{(with)}\;\;\;\;Z_{E}$ $:=\int\mathcal{D}\phi\;e^{-S_{E}[\phi]},$ (52)
where $S_{E}:=S_{E}^{(2)}+S_{E,\mathrm{int}}$ and $\mathcal{D}\phi$ is the (formal) functional measure. In perturbation theory, (52) is realized diagrammatically by Euclidean Feynman integrals with regulated propagators and vertices, which are absolutely convergent in the ultraviolet under the Euclidean damping assumptions on $F$.
The OS axioms are conditions on the family $\{S_{n}\}_{n\geq 0}$ of Euclidean $n$-point distributions that are necessary and sufficient for the existence of a corresponding relativistic QFT in the Wightman framework, and hence a positive Hilbert space, a vacuum, and unitary Poincaré symmetry OS73 ; OS75 ; GJ81 .
To state them, let $\mathcal{S}(\mathbb{R}^{4n})$ denote the Schwartz test functions on $\mathbb{R}^{4n}$ and $\mathcal{S}^{\prime}(\mathbb{R}^{4n})$ the tempered distributions. For $f\in\mathcal{S}(\mathbb{R}^{4n})$ we write:
$S_{n}(f):=\int_{\mathbb{R}^{4n}}d^{4n}x_{E}\;S_{n}(x_{E,1},\ldots,x_{E,n})\,f(x_{E,1},\ldots,x_{E,n})$ (53)
whenever this pairing is well-defined, distributional pairing.
(OS0) Regularity / temperedness.
For each $n\geq 0$, $S_{n}\in\mathcal{S}^{\prime}(\mathbb{R}^{4n})$, i.e. each $S_{n}$ is a tempered distribution.
(OS1) Euclidean invariance.
Let $E(4)=\mathbb{R}^{4}\rtimes O(4)$ be the Euclidean group acting on $\mathbb{R}^{4}$ by $x_{E}\mapsto Rx_{E}+a$ with $R\in O(4)$ and $a\in\mathbb{R}^{4}$. Then:
$S_{n}(Rx_{E,1}+a,\ldots,Rx_{E,n}+a)\;=\;S_{n}(x_{E,1},\ldots,x_{E,n})$ (54)
as distributions, for all $n\geq 0$.
(OS2) Symmetry.
For bosonic fields, $S_{n}$ is invariant under permutations of its arguments, so for any permutation $\pi\in S_{n}$:
$S_{n}(x_{E,\pi(1)},\ldots,x_{E,\pi(n)})\;=\;S_{n}(x_{E,1},\ldots,x_{E,n})$ (55)
as distributions.
(OS3) Reflection positivity.
Let $\theta:\mathbb{R}^{4}\to\mathbb{R}^{4}$ be reflection in Euclidean time:
$\theta(\tau,\mathbf{x}):=(-\tau,\mathbf{x}),$ (56)
and define the positive-time half-space $\mathbb{R}^{4}_{+}:=\{(\tau,\mathbf{x})\in\mathbb{R}^{4}:\tau>0\}$. Let $\mathcal{S}(\mathbb{R}^{4n}_{+})$ denote test functions on $\mathbb{R}^{4n}$ with support contained in $(\mathbb{R}^{4}_{+})^{n}$. For $f\in\mathcal{S}(\mathbb{R}^{4n})$ define the reflected-conjugated test function:
$(\Theta f)(x_{E,1},\ldots,x_{E,n}):=\overline{f(\theta x_{E,n},\ldots,\theta x_{E,1})}.$ (57)
Reflection positivity is the requirement that for every finite family $f_{k}\in\mathcal{S}(\mathbb{R}^{4n_{k}}_{+})$ and coefficients $c_{k}\in\mathbb{C}$:
$\sum_{k,\ell}\overline{c_{k}}\,c_{\ell}\;S_{n_{k}+n_{\ell}}(\Theta f_{k}\otimes f_{\ell})\;\geq\;0.$ (58)
(OS4) Cluster property.
Let $a\in\mathbb{R}^{4}$ and denote translation by $a$ as $x_{E}\mapsto x_{E}+a$. For separated clusters [^1] Clusters mean two or more groups of operator insertions in an $n$-point function that you separate far apart in Euclidean space(-time). Our quasi-locality affects how fast correlations die with separation, exponential vs power-law tails, etc., but clusters still just means those separated groups of insertions. The OS axiom is about the limit and factorization, not necessarily the detailed rate. one requires that for all $n,m\geq 0$:
$\lim_{\lambda\to+\infty}S_{n+m}(x_{E,1},\ldots,x_{E,n},\,y_{E,1}+\lambda a,\ldots,y_{E,m}+\lambda a)$ (59)
$=S_{n}(x_{E,1},\ldots,x_{E,n})\,S_{m}(y_{E,1},\ldots,y_{E,m}),$ (60)
in the sense of distributions, for any fixed nonzero $a$.
Theorem (Osterwalder–Schrader).
If the Schwinger functions $\{S_{n}\}_{n\geq 0}$ satisfy (OS0)–(OS4), then there exist: (i) a Hilbert space $\mathcal{H}$ with inner product $\langle\cdot,\cdot\rangle$, (ii) a cyclic vacuum vector $\Omega\in\mathcal{H}$ with $\|\Omega\|=1$, (iii) operator-valued tempered distributions (Wightman fields) on Minkowski space, and (iv) a strongly continuous unitary representation $U$ of the Poincaré group on $\mathcal{H}$ with spectrum condition, such that the corresponding Minkowski $n$-point Wightman distributions $W_{n}(x_{1},\ldots,x_{n}):=\langle\Omega,\phi(x_{1})\cdots\phi(x_{n})\Omega\rangle$ are obtained as boundary values of analytic continuations of the Euclidean functions $S_{n}$ under the Wick rotation $\tau=it$ with appropriate ordering prescriptions. Moreover, the reconstructed Wightman theory satisfies the Wightman axioms, in particular, positivity and Poincaré covariance, and the reconstruction is unique up to unitary equivalence OS73 ; OS75 ; GJ81 .
We emphasize what this means operationally that the Euclidean theory is not merely a calculational device; when (OS0)–(OS4) hold, it is the QFT, and Lorentzian correlators are obtained by analytic continuation from the Euclidean domain.
The OS framework cleanly separates two issues that are often conflated, existence and analyticity of Euclidean correlators versus reflection positivity needed for Hilbert space reconstruction.
Under Euclidean UV damping assumptions on $F$, Euclidean loop integrals defining the regulated Schwinger functions are absolutely convergent in the ultraviolet, and the Euclidean-first prescription defines analytic functions of external invariants away from the usual pinch singularities. In this sense, our Euclidean-first definition is aligned with the OS strategy: define $S_{n}$ in Euclidean signature first, and then continue.
Reflection positivity is not implied by mere UV finiteness or by the absence of additional poles. Even in the free scalar case, OS3 imposes a stringent constraint on the Euclidean two-point function NeebRP ; JaffeRP . For a translation-invariant Gaussian measure, the covariance $C_{E}(x_{E}-y_{E})$ with Fourier transform:
$\widetilde{C}_{E}(p_{E})=\frac{F(-p_{E}^{2}/M_{\ast}^{2})}{p_{E}^{2}+m^{2}},\qquad p_{E}^{2}:=(p_{E}^{0})^{2}+\mathbf{p}_{E}^{2},$ (61)
is reflection positive only for a restricted class of $\widetilde{C}_{E}$, equivalently, a restricted class of regulators that admit a positive Källén–Lehmann/Stieltjes-type representation NeebRP ; GJ81 . Thus, if one wants an OS-reconstructible Euclidean formulation for the fundamental fields, (OS3) becomes an admissibility criterion on the regulator beyond (A1)–(A2).
The Paley–Wiener analysis shows that demanding an entire UV-damping form factor forces the position-space kernel to have exponentially small but nonzero tails, so strict compact support is impossible. This is precisely the quasi-local regime where locality is recovered only in the limit $M_{\ast}\to\infty$. Since the classical OS theorem reconstructs a local Wightman theory, one should interpret it here in one of two logically distinct ways, either as a rigorous underpinning of the local limit $M_{\ast}\to\infty$, or as a benchmark set of axioms that can be imposed on appropriate gauge-invariant, or otherwise physical Euclidean observables, while allowing the regulated fundamental fields to be quasi-local at scales $\lesssim M_{\ast}^{-1}$.
The OS axioms articulate the precise additional conditions required to elevate the Euclidean-first prescription from UV-finite analytic continuation to a full Hilbert-space reconstruction. In our framework, UV finiteness and analytic continuation follow from the entire-function damping assumptions, while OS reflection positivity represents an additional, model-dependent constraint whose status must be checked, or imposed separately, especially in the presence of quasi-locality.
IX Discussion and outlook